OpenStudy (anonymous):

Write an equation of the line perpendicular to the line given by 2x + 3y = 7 and having the same y-intercept as the line given by 5x + 2y = 10. please show work

hartnn (hartnn):

can u find the y-intercept of the line 5x+2y=10 ??

OpenStudy (anonymous):

Y=3/2x+5 this is the answer i got which was right(at least the teacher said it was) but i have a problem showing the propper steps to come up with it. i have my way which is easy for me but is not always the right way. so i guess i am just looking to see the propper steps to coming up with the answer

hartnn (hartnn):

okk,i assume the equation of line in slope point form as y=mx+c i need to find slope m and y-intercept c since y-intercepts are equal for both lines,first i find the y-intercept of the line 5x+2y=10....which i find by putting x=0,so i get y=5 so my c=5 now to find m,i need slope of line 2x+3y = 7 because product of slopes of perpendicular lines is -1.......so slope of that line i get it as -2/3.....so m*(-2/3)=-1 so m=3/2....now i have both,m=3/2 and c= 5,i put it in y=mx+c to get y=3/2x+5 which u also got .. :) ok?ask if u have doubt in any of the steps...

OpenStudy (anonymous):

thank you :)